# Measuring cosmic baryon density with FRB and GW data

### v1

@@ Appendix B FRB data
FRB 20121102A 0.19273 557 188.4 82.9946 $33.1479$ 1 Chatterjee et al. (2017) Tendulkar et al. (2017)
FRB 20180301A 0.3304 552 151.7 93.2268 $4.6711$ 1 Bhandari et al. (2022)
FRB 20180924B 0.3214 361.42 40.5 326.1053 $-40.9$ 3 Bannister et al. (2019)
FRB 20181112A 0.4755 589.27 41.7 327.3485 $-52.9709$ 3 Prochaska et al. (2019)
FRB 20181223C 0.03024 111.61 19.9 180.9207 $27.5476$ 3 Bhardwaj et al. (2024)
FRB 20190102C 0.2912 364.5 57.4 322.4157 $-79.4757$ 3 Bhandari et al. (2020)
FRB 20190110C 0.12244 221.6 37.1 249.3185 $41.4434$ 2 Ibik et al. (2024)
FRB 20190303A 0.064 223.2 29.8 207.9958 $48.1211$ 2 Michilli et al. (2023)
FRB 20190418A 0.07132 182.78 70.2 65.8123 $16.0738$ 3 Bhardwaj et al. (2024)
FRB 20190523A 0.66 760.8 37.2 207.065 $72.4697$ 3 Ravi et al. (2019)
FRB 20190608B 0.11778 338.7 37.3 334.0199 $-7.8983$ 3 Chittidi et al. (2021)
FRB 20190611B 0.3778 321.4 57.8 320.7456 $-79.3976$ 3 Heintz et al. (2020)
FRB 20190614D 0.6 959.2 87.8 65.0755 $73.7067$ 3 Law et al. (2020)
FRB 20190711A 0.522 593.1 56.5 329.4193 $-80.358$ 1 Heintz et al. (2020)
FRB 20190714A 0.2365 504.13 38.5 183.9797 $-13.021$ 3 Heintz et al. (2020)
FRB 20191001A 0.234 506.92 44.2 323.3513 $-54.7478$ 3 Heintz et al. (2020)
FRB 20191106C 0.10775 332.2 25 199.5801 $42.9997$ 2 Ibik et al. (2024)
FRB 20191228A 0.2432 297.5 32.9 344.4304 $-28.5941$ 3 Bhandari et al. (2022)
FRB 20200223B 0.06024 201.8 45.6 8.2695 $8.83132$ Ibik et al. (2024)
FRB 20200430A 0.1608 380.1 27.2 229.7064 $12.3763$ 3 Heintz et al. (2020)
FRB 20200723B 0.0085 244 33 190.1583 $-5.135$ 3 Shin et al. (2024)
FRB 20200906A 0.3688 577.8 35.8 53.4962 $-14.0832$ 3 Bhandari et al. (2022)
FRB 20201124A 0.098 413.52 139.9 77.0146 $6.06072$ Ravi et al. (2022)
FRB 20210117A 0.2145 729.1 34.4 339.9792 $-16.1515$ 3 Bhandari et al. (2023)
FRB 20210320C 0.2797 384.8 39.3 204.4608 $-16.1227$ 3 Shannon et al. (2025)
FRB 20210410D 0.1415 571.2 56.2 326.0863 $-79.3182$ 3 Caleb et al. (2023)
FRB 20210603A 0.1772 500.147 39.5 10.2741 $21.2263$ 3 Cassanelli et al. (2024)
FRB 20210807D 0.1293 251.3 121.2 299.2214 $-0.7624$ 3 Gordon et al. (2023)
FRB 20211127I 0.0469 234.83 42.5 199.8082 $-18.8378$ 3 Gordon et al. (2023)
FRB 20211203C 0.3439 636.2 63.7 204.5625 $-31.3801$ 3 Gordon et al. (2023)
FRB 20211212A 0.0715 206 27.1 157.3509 $1.676889$ 3 Gordon et al. (2023)
FRB 20220105A 0.2785 583 22 208.8039 $22.4665$ 3 Gordon et al. (2023)
FRB 20220204A 0.4012 612.584 50.7 274.2263 $69.7225$ 3 Sharma et al. (2024)
FRB 20220207C 0.04304 262.38 76.1 310.1995 $72.8823$ 3 Law et al. (2024)
FRB 20220208A 0.351 437 101.6 319.3483 $71.54$ 3 Sharma et al. (2024)
FRB 20220307B 0.2507 499.15 128.2 350.8745 $72.1924$ 3 Law et al. (2024)
FRB 20220310F 0.477958 462.24 46.3 134.7204 $73.4908$ 3 Law et al. (2024)
FRB 20220330D 0.3714 468.1 38.6 165.7256 $71.7535$ 3 Sharma et al. (2024)
FRB 20220418A 0.622 623.25 36.7 219.1056 $70.0959$ 3 Law et al. (2024)
FRB 20220501C 0.381 449.5 30.6 352.3792 $-32.4907$ 3 Shannon et al. (2025)
FRB 20220506D 0.30039 396.97 84.6 318.0448 $72.8273$ 3 Law et al. (2024)
FRB 20220509G 0.0894 269.53 55.6 282.67 $70.2438$ 3 Law et al. (2024)
FRB 20220521B 1.354 1342.9 138.8 351.036 $71.138$ 3 Shannon et al. (2025)
FRB 20220529A 0.1839 246 40 19.1042 $20.6325$ 1 Li et al. (2025)
FRB 20220610A 1.016 1458.15 31 351.0732 $-33.5137$ 3 Ryder et al. (2023)
FRB 20220717A 0.36295 637.34 118.3 293.3042 $-19.2877$ 3 Rajwade et al. (2024)
FRB 20220725A 0.1926 290.4 30.7 353.3152 $-35.9902$ 3 Shannon et al. (2025)
FRB 20220726A 0.361 686.55 89.5 73.94567 $69.9291$ 3 Sharma et al. (2024)
FRB 20220825A 0.241397 651.24 78.5 311.9815 $72.585$ 3 Law et al. (2024)
FRB 20220914A 0.1139 631.28 54.7 282.0568 $73.3369$ 3 Law et al. (2024)
FRB 20220918A 0.491 656.8 153.1 17.5921 $70.8113$ 3 Shannon et al. (2025)
FRB 20220920A 0.158239 314.99 39.9 240.2571 $70.9188$ 3 Law et al. (2024)
FRB 20221012A 0.284669 441.08 54.3 280.7987 $70.5242$ 3 Law et al. (2024)
FRB 20221027A 0.229 452.5 47.2 129.6104 $71.7315$ 3 Sharma et al. (2024)
FRB 20221029A 0.975 1391.75 43.8 141.9634 $72.4523$ 3 Sharma et al. (2024)
FRB 20221101B 0.2395 491.554 131.2 342.2162 $70.6812$ 3 Sharma et al. (2024)
FRB 20221106A 0.2044 343.8 34.8 56.7048 $-25.5698$ 3 Shannon et al. (2025)
FRB 20221113A 0.2505 411.027 91.7 71.411 $70.3074$ 3 Sharma et al. (2024)
FRB 20221116A 0.2764 643.448 132.3 21.2102 $72.6539$ 3 Sharma et al. (2024)
FRB 20221219A 0.553 706.708 44.4 257.6298 $71.6268$ 3 Sharma et al. (2024)
FRB 20230124A 0.0939 590.574 38.6 231.9163 $70.9681$ 3 Sharma et al. (2024)
FRB 20230203A 0.1464 420.1 67.299 151.66159 $5.69413$ Amiri et al. (2025)
FRB 20230216A 0.531 828 38.5 155.9717 $1.4678$ 3 Sharma et al. (2024)
FRB 20230222A 0.1223 706.1 33.31 106.96036 $11.22452$ 3 Amiri et al. (2025)
FRB 20230222B 0.11 187.8 56.9977 238.73912 $0.89873$ Amiri et al. (2025)
FRB 20230307A 0.271 608.9 37.6 177.7813 $71.6956$ 3 Sharma et al. (2024)
FRB 20230311A 0.1918 364.3 67.2352 91.10966 $55.94595$ 3 Amiri et al. (2025)
FRB 20230501A 0.301 532.5 125.6 340.0272 $70.9222$ 3 Sharma et al. (2024)
FRB 20230506C 0.3896 772 68 12.0998 $42.0061$ 2 Anna-Thomas et al. (2025)
FRB 20230526A 0.157 361.4 31.9 22.2326 $-52.7173$ 3 Shannon et al. (2025)
FRB 20230626A 0.327 451.2 39.2 240.7125 $71.7142$ 3 Sharma et al. (2024)
FRB 20230628A 0.1265 344.952 39 166.7867 $72.2818$ 3 Sharma et al. (2024)
FRB 20230703A 0.1184 291.3 38.1566 184.62445 $48.72993$ 3 Amiri et al. (2025)
FRB 20230708A 0.105 411.51 60.3 303.1155 $-55.3563$ 3 Shannon et al. (2025)
FRB 20230712A 0.4525 587.567 39.2 167.3585 $72.5578$ 3 Sharma et al. (2024)
FRB 20230730A 0.2115 312.5 114.6465 54.66456 $3.15933$ Amiri et al. (2025)
FRB 20230808F 0.3472 653.2 31 53.3041 $-51.9353$ 3 Hanmer et al. (2025)
FRB 20230814A 0.553 696.4 104.8 335.9748 $73.0259$ 3 Connor et al. (2025)
FRB 20230902A 0.3619 440.1 34.1 52.1398 $-47.3335$ 3 Shannon et al. (2025)
FRB 20230930A 0.0925 456 70 10.5 $41.4$ 3 Anna-Thomas et al. (2025)
FRB 20231011A 0.0783 186.3 58.5524 18.2411 $41.7491$ 3 Amiri et al. (2025)
FRB 20231017A 0.245 344.2 31.1515 346.75429 $6.652683$ Amiri et al. (2025)
FRB 20231025B 0.3238 368.7 65.5787 270.78807 $63.98908$ 3 Amiri et al. (2025)
FRB 20231120A 0.07 438.9 43.8 143.6169 $71.7574$ 3 Sharma et al. (2024)
FRB 20231123A 0.0729 302.1 37.1085 82.62325 $4.47554$ 3 Amiri et al. (2025)
FRB 20231123B 0.2625 396.7 40.2 242.5382 $70.7851$ 3 Sharma et al. (2024)
FRB 20231128A 0.1079 331.6 64.7368 199.5782 $42.99271$ 1 Amiri et al. (2025)
FRB 20231204A 0.0644 221 34.9372 207.99903 $48.116$ 2 Amiri et al. (2025)
FRB 20231206A 0.0659 457.7 37.5114 112.44284 $56.25627$ 3 Amiri et al. (2025)
FRB 20231220A 0.3355 491.2 49.9 123.9087 $73.6599$ 3 Connor et al. (2025)
FRB 20231226A 0.1569 329.9 38.1 155.3638 $6.1103$ 3 Shannon et al. (2025)
FRB 20231229A 0.019 198.5 65.1335 26.46783 $5.112923$ Amiri et al. (2025)
FRB 20231230A 0.0298 131.4 32.2305 72.79761 $2.39398$ 3 Amiri et al. (2025)
FRB 20240114A 0.13 527.65 49.7 321.9161 $4.3292$ 1 Tian et al. (2024)
FRB 20240119A 0.376 483.1 38 224.4672 $71.6118$ 3 Connor et al. (2025)
FRB 20240123A 0.968 1462 90.2 68.2625 $71.9453$ 3 Connor et al. (2025)
FRB 20240201A 0.042729 374.5 38.6 149.9056 $14.088$ 3 Shannon et al. (2025)
FRB 20240209A 0.1384 176.518 63.1 289.8504 $86.0609$ 2 Eftekhari et al. (2025)
FRB 20240210A 0.023686 283.73 28.7 8.7796 $-28.2708$ 3 Shannon et al. (2025)
FRB 20240213A 0.1185 357.4 40 166.1683 $74.0754$ 3 Connor et al. (2025)
FRB 20240215A 0.21 549.5 47.9 268.4413 $70.2324$ 3 Connor et al. (2025)
FRB 20240229A 0.287 491.15 38 169.9835 $70.6762$ 3 Connor et al. (2025)
FRB 20240304A 0.2423 652.6 73.9 136.3304 $-16.2282$ 3 Shannon et al. (2025)
FRB 20240310A 0.127 601.8 30.1 17.6219 $-44.4394$ 3 Shannon et al. (2025)
Note. — The results are derived from $104$ localized FRBs alone or in combination with $47$ GW events, under different assumptions. For all cases, the $1\sigma$ constraints on DM parameters are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$. $H_{0}$ is in units of km s-1  Mpc-1.

### v2

@@ Appendix B FRB data
FRB 20121102A 0.19273 557 188.4 82.9946 $33.1479$ 1 Chatterjee:2017dqg; Tendulkar2016
FRB 20180301A 0.3304 552 151.7 93.2268 $4.6711$ 1 Bhandari2022
FRB 20180924B 0.3214 361.42 40.5 326.1053 40.9-40.9 3 Bannister2019
FRB 20181112A 0.4755 589.27 41.7 327.3485 52.9709-52.9709 3 Prochaska2019
FRB 20181223C 0.03024 111.61 19.9 180.9207 $27.5476$ 3 Bhardwaj2024
FRB 20190102C 0.2912 364.5 57.4 322.4157 79.4757-79.4757 3 Bhandari2020
FRB 20190110C 0.12244 221.6 37.1 249.3185 $41.4434$ 2 Ibik2024
FRB 20190303A 0.064 223.2 29.8 207.9958 $48.1211$ 2 Michilli2023
FRB 20190418A 0.07132 182.78 70.2 65.8123 $16.0738$ 3 Bhardwaj2024
FRB 20190523A 0.66 760.8 37.2 207.065 $72.4697$ 3 Ravi2019
FRB 20190608B 0.11778 338.7 37.3 334.0199 7.8983-7.8983 3 Chittidi2021
FRB 20190611B 0.3778 321.4 57.8 320.7456 79.3976-79.3976 3 Heintz2020
FRB 20190614D 0.6 959.2 87.8 65.0755 $73.7067$ 3 Law2020
FRB 20190711A 0.522 593.1 56.5 329.4193 80.358-80.358 1 Heintz2020
FRB 20190714A 0.2365 504.13 38.5 183.9797 13.021-13.021 3 Heintz2020
FRB 20191001A 0.234 506.92 44.2 323.3513 54.7478-54.7478 3 Heintz2020
FRB 20191106C 0.10775 332.2 25 199.5801 $42.9997$ 2 Ibik2024
FRB 20191228A 0.2432 297.5 32.9 344.4304 28.5941-28.5941 3 Bhandari2022
FRB 20200223B 0.06024 201.8 45.6 8.2695 $28.8313$ 2 Ibik2024
FRB 20200430A 0.1608 380.1 27.2 229.7064 $12.3763$ 3 Heintz2020
FRB 20200723B 0.0085 244 33 190.1583 5.135-5.135 3 2024arXiv241007307S
FRB 20200906A 0.3688 577.8 35.8 53.4962 14.0832-14.0832 3 Bhandari2022
FRB 20201124A 0.098 413.52 139.9 77.0146 $26.0607$ 2 Ravi2022
FRB 20210117A 0.2145 729.1 34.4 339.9792 16.1515-16.1515 3 Bhandari2023
FRB 20210320C 0.2797 384.8 39.3 204.4608 16.1227-16.1227 3 Shannon2024
FRB 20210410D 0.1415 571.2 56.2 326.0863 79.3182-79.3182 3 Caleb2023
FRB 20210603A 0.1772 500.147 39.5 10.2741 $21.2263$ 3 Cassanelli2023
FRB 20210807D 0.1293 251.3 121.2 299.2214 0.7624-0.7624 3 Gordon2023
FRB 20211127I 0.0469 234.83 42.5 199.8082 18.8378-18.8378 3 Gordon2023
FRB 20211203C 0.3439 636.2 63.7 204.5625 31.3801-31.3801 3 Gordon2023
FRB 20211212A 0.0715 206 27.1 157.3509 $1.676889$ 3 Gordon2023
FRB 20220105A 0.2785 583 22 208.8039 $22.4665$ 3 Gordon2023
FRB 20220204A 0.4012 612.584 50.7 274.2263 $69.7225$ 3 Sharma2024
FRB 20220207C 0.04304 262.38 76.1 310.1995 $72.8823$ 3 Law2024
FRB 20220208A 0.351 437 101.6 319.3483 $71.54$ 3 Sharma2024
FRB 20220307B 0.2507 499.15 128.2 350.8745 $72.1924$ 3 Law2024
FRB 20220310F 0.477958 462.24 46.3 134.7204 $73.4908$ 3 Law2024
FRB 20220330D 0.3714 468.1 38.6 165.7256 $71.7535$ 3 Sharma2024
FRB 20220418A 0.622 623.25 36.7 219.1056 $70.0959$ 3 Law2024
FRB 20220501C 0.381 449.5 30.6 352.3792 32.4907-32.4907 3 Shannon2024
FRB 20220506D 0.30039 396.97 84.6 318.0448 $72.8273$ 3 Law2024
FRB 20220509G 0.0894 269.53 55.6 282.67 $70.2438$ 3 Law2024
FRB 20220521B 1.354 1342.9 138.8 351.036 $71.138$ 3 Shannon2024
FRB 20220529A 0.1839 246 40 19.1042 $20.6325$ 1 Li:2025ckl
FRB 20220610A 1.016 1458.15 31 351.0732 33.5137-33.5137 3 Ryder2023
FRB 20220717A 0.36295 637.34 118.3 293.3042 19.2877-19.2877 3 Rajwade2024
FRB 20220725A 0.1926 290.4 30.7 353.3152 35.9902-35.9902 3 Shannon2024
FRB 20220726A 0.361 686.55 89.5 73.94567 $69.9291$ 3 Sharma2024
FRB 20220825A 0.241397 651.24 78.5 311.9815 $72.585$ 3 Law2024
FRB 20220914A 0.1139 631.28 54.7 282.0568 $73.3369$ 3 Law2024
FRB 20220918A 0.491 656.8 153.1 17.5921 $70.8113$ 3 Shannon2024
FRB 20220920A 0.158239 314.99 39.9 240.2571 $70.9188$ 3 Law2024
FRB 20221012A 0.284669 441.08 54.3 280.7987 $70.5242$ 3 Law2024
FRB 20221027A 0.229 452.5 47.2 129.6104 $71.7315$ 3 Sharma2024
FRB 20221029A 0.975 1391.75 43.8 141.9634 $72.4523$ 3 Sharma2024
FRB 20221101B 0.2395 491.554 131.2 342.2162 $70.6812$ 3 Sharma2024
FRB 20221106A 0.2044 343.8 34.8 56.7048 25.5698-25.5698 3 Shannon2024
FRB 20221113A 0.2505 411.027 91.7 71.411 $70.3074$ 3 Sharma2024
FRB 20221116A 0.2764 643.448 132.3 21.2102 $72.6539$ 3 Sharma2024
FRB 20221219A 0.553 706.708 44.4 257.6298 $71.6268$ 3 Sharma2024
FRB 20230124A 0.0939 590.574 38.6 231.9163 $70.9681$ 3 Sharma2024
FRB 20230203A 0.1464 420.1 67.299 151.66159 $35.6941$ 3 Amiri:2025sbi
FRB 20230216A 0.531 828 38.5 155.9717 $1.4678$ 3 Sharma2024
FRB 20230222A 0.1223 706.1 33.31 106.96036 $11.22452$ 3 Amiri:2025sbi
FRB 20230222B 0.11 187.8 56.9977 238.73912 $30.8987$ 3 Amiri:2025sbi
FRB 20230307A 0.271 608.9 37.6 177.7813 $71.6956$ 3 Sharma2024
FRB 20230311A 0.1918 364.3 67.2352 91.10966 $55.94595$ 3 Amiri:2025sbi
FRB 20230501A 0.301 532.5 125.6 340.0272 $70.9222$ 3 Sharma2024
FRB 20230506C 0.3896 772 68 12.0998 $42.0061$ 2 AnnaThomas2025
FRB 20230526A 0.157 361.4 31.9 22.2326 52.7173-52.7173 3 Shannon2024
FRB 20230626A 0.327 451.2 39.2 240.7125 $71.7142$ 3 Sharma2024
FRB 20230628A 0.1265 344.952 39 166.7867 $72.2818$ 3 Sharma2024
FRB 20230703A 0.1184 291.3 38.1566 184.62445 $48.72993$ 3 Amiri:2025sbi
FRB 20230708A 0.105 411.51 60.3 303.1155 55.3563-55.3563 3 Shannon2024
FRB 20230712A 0.4525 587.567 39.2 167.3585 $72.5578$ 3 Sharma2024
FRB 20230730A 0.2115 312.5 114.6465 54.66456 $33.1593$ 3 Amiri:2025sbi
FRB 20230808F 0.3472 653.2 31 53.3041 51.9353-51.9353 3 2025MNRAS.538.1800H
FRB 20230814A 0.553 696.4 104.8 335.9748 $73.0259$ 3 connor2025gas
FRB 20230902A 0.3619 440.1 34.1 52.1398 47.3335-47.3335 3 Shannon2024
FRB 20230930A 0.0925 456 70 10.5 $41.4$ 3 AnnaThomas2025
FRB 20231011A 0.0783 186.3 58.5524 18.2411 $41.7491$ 3 Amiri:2025sbi
FRB 20231017A 0.245 344.2 31.1515 346.75429 $36.65268$ 3 Amiri:2025sbi
FRB 20231025B 0.3238 368.7 65.5787 270.78807 $63.98908$ 3 Amiri:2025sbi
FRB 20231120A 0.07 438.9 43.8 143.6169 $71.7574$ 3 Sharma2024
FRB 20231123A 0.0729 302.1 37.1085 82.62325 $4.47554$ 3 Amiri:2025sbi
FRB 20231123B 0.2625 396.7 40.2 242.5382 $70.7851$ 3 Sharma2024
FRB 20231128A 0.1079 331.6 64.7368 199.5782 $42.99271$ 1 Amiri:2025sbi
FRB 20231204A 0.0644 221 34.9372 207.99903 $48.116$ 2 Amiri:2025sbi
FRB 20231206A 0.0659 457.7 37.5114 112.44284 $56.25627$ 3 Amiri:2025sbi
FRB 20231220A 0.3355 491.2 49.9 123.9087 $73.6599$ 3 connor2025gas
FRB 20231226A 0.1569 329.9 38.1 155.3638 $6.1103$ 3 Shannon2024
FRB 20231229A 0.019 198.5 65.1335 26.46783 $35.11292$ 3 Amiri:2025sbi
FRB 20231230A 0.0298 131.4 32.2305 72.79761 $2.39398$ 3 Amiri:2025sbi
FRB 20240114A 0.13 527.65 49.7 321.9161 $4.3292$ 1 Tian:2024ygd
FRB 20240119A 0.376 483.1 38 224.4672 $71.6118$ 3 connor2025gas
FRB 20240123A 0.968 1462 90.2 68.2625 $71.9453$ 3 connor2025gas
FRB 20240201A 0.042729 374.5 38.6 149.9056 $14.088$ 3 Shannon2024
FRB 20240209A 0.1384 176.518 63.1 289.8504 $86.0609$ 2 2025ApJ...979L..22E
FRB 20240210A 0.023686 283.73 28.7 8.7796 28.2708-28.2708 3 Shannon2024
FRB 20240213A 0.1185 357.4 40 166.1683 $74.0754$ 3 connor2025gas
FRB 20240215A 0.21 549.5 47.9 268.4413 $70.2324$ 3 connor2025gas
FRB 20240229A 0.287 491.15 38 169.9835 $70.6762$ 3 connor2025gas
FRB 20240304A 0.2423 652.6 73.9 136.3304 16.2282-16.2282 3 Shannon2024
FRB 20240310A 0.127 601.8 30.1 17.6219 44.4394-44.4394 3 Shannon2024
Table 2: Continued